OpenStudy (superhelp101):

You run a canoe-rental business on a small river in Ohio. You currently charge $12 per canoe and average 36 rentals a day. An industry journal says that, for every fifty-cent increase in rental price, the average business can expect to lose two rentals a day. Use this information to attempt to maximize your income. What should you charge?

OpenStudy (superhelp101):

@aum @ganeshie8

ganeshie8 (ganeshie8):

say : n = number of rentals p = chanrge per canoe

ganeshie8 (ganeshie8):

income can be expressed as : \[\large \text{income} = n\times p\] yes ?

OpenStudy (superhelp101):

yes

ganeshie8 (ganeshie8):

good, say change in price of canoe = x dollars then, change in rental would be : 4x right ?

OpenStudy (superhelp101):

yes

ganeshie8 (ganeshie8):

nope

OpenStudy (superhelp101):

oops

ganeshie8 (ganeshie8):

good, say change in price of canoe = x dollars then, change in number of rentals would be : \(\large \color{Red}{-}\)4x right ?

OpenStudy (superhelp101):

yes

ganeshie8 (ganeshie8):

it has to be negative because the number of rentals decrease if u increase the price

ganeshie8 (ganeshie8):

so can we express the income as below : \[\large \text{income} = n\times p\] \[\large \text{income} = (36-4x)\times (12+x)\] ?

OpenStudy (superhelp101):

okay I understand how we would then solve it, but don't get how we formed the equation

OpenStudy (superhelp101):

like how did we get -4x

ganeshie8 (ganeshie8):

`An industry journal says that, for every fifty-cent increase in rental price, the average business can expect to lose two rentals a day. `

ganeshie8 (ganeshie8):

that says, every 50cent increase in price lowers the number of rentals by 2

ganeshie8 (ganeshie8):

that means every 1 dollar increase in price lowers the number of rentals by 4 yes ?

OpenStudy (superhelp101):

okay

OpenStudy (superhelp101):

actually how do you solve it after ?